# Generative thermodynamic computing

### v2

@@ Abstract
In this mode of operation, the thermodynamic computer resembles a nonequilibrium, continuous-spin analog of a Boltzmann machine, a statistical mechanical model that represents probability distributions over binary variables Hinton (2017); Salakhutdinov and Hinton (2009). Given that a Boltzmann machine encodes information in its equilibrium or Boltzmann distribution, we refer to the device considered here, which uses Langevin trajectories to perform a calculation, as a Langevin computer.
@@
The first sum in Eq. (2), which runs over the $N$ units, sets the intrinsic couplings of the computer. For $J_{2}>0,J_{4}=0$ we have a linear model Aifer et al. (2024), whose unit activations are linear functions of their inputs, while for $J_{2},J_{4}>0$  [^1] Note that the choice $J_{2}<0,J_{4}>0$ creates bistable units, analogous to the p-spins in the field of probabilistic computing Kaiser and Datta (2021). The choice $J_{2}>0$ creates units with one stable state, analogous to the s-units of thermodynamic computing Melanson et al. (2025). we have a nonlinear model that can act as the thermodynamic analog of a neural network Whitelam and Casert (2024). The remaining terms in (2) contain the trainable parameters of the computer. The parameters $b_{i}$ are input signals or biases applied to each unit. The parameters $J_{ij}$ are pairwise couplings between units, inspired by the bilinear interactions of the thermodynamic computers of Refs. Aifer et al. (2024); Melanson et al. (2025), with the sum running over all inter-unit connections. Eq. (2) describes a thermodynamic computer of arbitrary connectivity, and the following discussion applies to the same.
@@
$-\ln P^{\rm step}_{\bm{\theta}}(\Delta{\bm{x}})=\sum_{i=1}^{N}\frac{\left(\Delta x_{i}+\mu\mskip 0.75mu\partial_{i}V_{\bm{\theta}}({\bm{x}})\Delta t\right)^{2}}{4\mu\mskip 0.75muk_{\rm B}T\Delta t},$ (6)
@@
$-\ln\tilde{P}^{\rm step}_{\bm{\theta}}(\Delta{\bm{x}})=\sum_{i=1}^{N}\frac{\left(-\Delta x_{i}+\mu\mskip 0.75mu\partial_{i}V_{\bm{\theta}}({\bm{x}}^{\prime})\Delta t\right)^{2}}{4\mu\mskip 0.75muk_{\rm B}T\Delta t}.$ (7)
@@
$-\frac{\partial}{\partial b_{i}}\ln\tilde{P}^{\rm step}_{\bm{\theta}}(\Delta{\bm{x}})=\frac{-\Delta x_{i}+\mu\mskip 0.75mu\partial_{i}V_{\bm{\theta}}({\bm{x}}^{\prime})\Delta t}{2k_{\rm B}T},$ (11)
@@
Training proceeds by minimizing the negative log-probability $-\ln P_{\bm{\theta}}[\tilde{\omega}]$. According to the fluctuation relation (14), and given that the reference process is fixed, this is equivalent to minimizing $-Q_{\bm{\theta}}(\omega)$, the negative total heat dissipated by the denoising computer when generating the noising trajectory $\omega$. Given that heat changes sign upon time reversal, the learning process therefore minimizes the heat $Q_{\bm{\theta}}(\tilde{\omega})=-Q_{\bm{\theta}}(\omega)$ emitted by the trained computer along the trajectory $\tilde{\omega}$, i.e. as it generates structure from noise. In this sense, the trained dynamics is thermodynamically optimal: it is the dynamics that reconstructs the imposed data with the least heat emitted or entropy produced.

### v3

@@ Abstract
In this mode of operation the thermodynamic computer resembles a nonequilibrium, continuous-spin analog of a Boltzmann machine, a statistical mechanical model that represents probability distributions over binary variables Hinton (2017); Salakhutdinov and Hinton (2009). The key difference is that a Boltzmann machine encodes information in its equilibrium Boltzmann distribution, whereas our device runs on a physical clock: the computation is carried out by the dynamics of the system at a designated time, with no requirement to attain equilibrium. We therefore refer to the device considered here, which uses Langevin trajectories to perform a calculation, as a Langevin computer.
@@
The first sum in Eq. (2), which runs over the $N$ units, sets the intrinsic couplings of the computer. For $J_{4}=0$ we have a linear model Aifer et al. (2024), whose unit activations are linear functions of their inputs, while for $J_{4}>0$, the case we consider, we have a nonlinear model that can act as the thermodynamic analog of a neural network Whitelam and Casert (2024) (positive $J_{4}$ also ensures the thermodynamic stability of the computer as the $J_{ij}$ are adjusted). We consider the case $J_{2}>0$, which creates units with one stable state, analogous to the s-units of thermodynamic computing Melanson et al. (2025). The alternative choice, $J_{2}<0$, creates bistable units, analogous to the p-spins in the field of probabilistic computing Kaiser and Datta (2021).
The remaining terms in (2) contain the trainable parameters of the computer. The parameters $b_{i}$ are input signals or biases applied to each unit. The parameters $J_{ij}$ are pairwise couplings between units, inspired by the bilinear interactions of the thermodynamic computers of Refs. Aifer et al. (2024); Melanson et al. (2025), with the sum running over all inter-unit connections. Eq. (2) describes a thermodynamic computer of arbitrary connectivity, and the following discussion applies to the same.
@@
$\ln P^{\rm step}_{\bm{\theta}}(\Delta{\bm{x}})=\sum_{i=1}^{N}\frac{\left(\Delta x_{i}+\mu\mskip 0.75mu\partial_{i}V_{\bm{\theta}}({\bm{x}})\Delta t\right)^{2}}{4\mu\mskip 0.75muk_{\rm B}T\Delta t},$ (6)
@@
$\ln\tilde{P}^{\rm step}_{\bm{\theta}}(\Delta{\bm{x}})=\sum_{i=1}^{N}\frac{\left(-\Delta x_{i}+\mu\mskip 0.75mu\partial_{i}V_{\bm{\theta}}({\bm{x}}^{\prime})\Delta t\right)^{2}}{4\mu\mskip 0.75muk_{\rm B}T\Delta t}.$ (7)
@@
$\frac{\partial}{\partial b_{i}}\ln\tilde{P}^{\rm step}_{\bm{\theta}}(\Delta{\bm{x}})=\frac{-\Delta x_{i}+\mu\mskip 0.75mu\partial_{i}V_{\bm{\theta}}({\bm{x}}^{\prime})\Delta t}{2k_{\rm B}T},$ (11)
@@
Training proceeds by minimizing the negative log-probability $\ln P_{\bm{\theta}}[\tilde{\omega}$]. According to the fluctuation relation (14), and given that the reference process is fixed, this is equivalent to minimizing $Q_{\bm{\theta}}(\omega)$, the negative total heat dissipated by the denoising computer when generating the noising trajectory $\omega$. Given that heat changes sign upon time reversal, the learning process therefore minimizes the heat $Q_{\bm{\theta}}(\tilde{\omega})=-Q_{\bm{\theta}}(\omega)$ emitted by the trained computer along the trajectory $\tilde{\omega}$, i.e. as it generates structure from noise. In this sense, the trained dynamics is thermodynamically optimal: it is the dynamics that reconstructs the imposed data with the least heat emitted or entropy produced.
Thermodynamic advantage — We can estimate the thermodynamic advantage, the ratio of the energy costs of digital to thermodynamic computation, by considering the energy scales of denoising using a digital neural network and a hardware version of our simulated thermodynamic computer. The basic energy scale of a digital neural network is set by a multiply-accumulate (MAC) operation, which differs by hardware implementation but is typically about 1 pJ Horowitz (2014), or $2.4\times 10^{8}\,k_{\rm B}T$ at room temperature. For a modest multilayer perceptron denoiser ($784\!\to\!128\!\to\!128\!\to\!784$), a single denoising step requires $\sim 2.2\times 10^{5}$ MACs. If we make the very conservative assumption that the denoiser is used only 10 times within a denoising trajectory, the order-of-magnitude energy budget of denoising using a neural network is not less than $5\times 10^{14}\,k_{\rm B}T$.
The energy cost of the thermodynamic computer is much smaller. We can calculate the heat emitted by the computer from the difference of the potential energy (2) between the start and end of a trajectory, $Q=V_{{\bm{\theta}}}({\bm{x}}(0))-V_{{\bm{\theta}}}({\bm{x}}(t_{\rm f}))$ (no work is done on the computer after the hidden-unit biases are established). Over $1000$ independent denoising trajectories of the trained computer we calculate a mean heat emission of $\langle Q\rangle=2.9\times 10^{3}\,k_{\rm B}T$, with standard deviation $3.5\times 10^{2}\,k_{\rm B}T$. Comparing this value to the digital estimate gives a ratio of more than $10^{11}$. That is, if implemented in hardware, the thermodynamic computer would be more than ten orders of magnitude more energy efficient than a digital neural network. The example shown here is rudimentary by the standards of state-of-the-art diffusion models, but shows the potential, and potential energy savings, of thermodynamic computation.